1gm ice mixed with 1gm steam find final temperature of mixture
1
Expert's answer
2020-10-07T07:27:18-0400

Heat required to melt 1g of ice,

"Q_1 = mL = 1 \times 80 = 80 cal"


Heat required by 1g of water to boil from "0^{\circ} C" at "100^{\circ} C" ,

"Q_2 = ms\Delta T = 1 \times1(100-0) = 100 cal"


Total heat required for 1g of ice to boil,

"Q = Q_1 +Q_2 = 80+100=180 cal"


Heat with 1g ice to condense 1g of water at "100^{\circ} C" ,

"Q' = mL = 1 \times 536 = 536 cal"


Since total steam will not be condensed hence final temperature of the system will be "100^{\circ} C"